OpenStudy (selena2345):

Sylvie's age is 5 years less than half Katie's age. If Sylvie is 11 years old, what is Katie's age ? A:8 Years old B: 12 years old C: 27 years old D: 32 years old

OpenStudy (anonymous):

S = 11 S = K/2 -5

OpenStudy (selena2345):

what ? @zimmah

OpenStudy (anonymous):

Sylvie = 11 Sylvie = Katy /2 -5 so Katy = Sylvie *2 +5 11*2+5

OpenStudy (selena2345):

Thank u @zimmah

OpenStudy (anonymous):

oh actually that last step is not correct

OpenStudy (anonymous):

it's +10 not +5 because you have to multiply the 5 by 2 as well